# Modeling error correction with Lindblad dynamics and approximate channels

### v1

@@ V.6 Nonsymmetric reference frame
However, when we are interested in the Pauli approximation, these two representations for the nonsymmetric reference frame are different. This stems from the fact that the Pauli approximation is done on the 1Q noise terms and 2Q noise terms separately. These differences can be seen in Fig. 7, for $\tilde{h}=2\pi\times 5\,$kHz, ${\zeta}=2\pi\times 3\,$kHz, $T_{1}=150$ $\mu$s and $T_{2}=100$ $\mu$s. In Fig. 7(a) we see that for the initial state $|+\rangle_{L}$ in the “2Q-adjusted” representation, the Pauli approximation begins to deviate already at $t=0.3$ $\mu$s, while in the “1Q-adjusted” representation, the Pauli approximation remains accurate up to longer times. The initial state $|0\rangle_{L}$, on the other hand is significantly different, in that the Pauli approximation does not perform well regardless of the representation, as can be seen in Fig. 7(b). In total, it seems that for these noise parameters, the 1Q-adjusted representation is somewhat beneficial for the Pauli approximation, though this may depend on the specific noise parameters. Again we see that the Pauli approximation is sensitive to the specific noise parameters.
@@ VI Summary and Outlook
The focus of this work is the importance of physically-motivated noise models when determining the success rate of an error correction code to recover logical states subject to noisy dynamics. The Lindblad model studied in this work includes the dominant errors in superconducting quantum computing systems [19], or can be modified in order to better model those  [^5] We note that the effect of frequency shifts caused by stochastic charge-parity fluctuation in superconducting qubits [19], can be absorbed into each qubit’s $h_{i}$ value and averaged over all possible frequencies., and could be relevant for other systems as well [56, 57, 58, 59, 60, 61]. It encompasses two-qubit terms and thus extends beyond single-qubit terms that are often assumed. Moreover, it incorporates noncommuting terms, thereby introducing nontrivial dynamics into the system.
@@
Overall, we have determined that for a Lindbladian of the form presented in Sec. II, the composite-channel approximation (constructed explicitly to account for 2Q noise) is a very good approximation for relevant noise parameters and the time scales that we study. This approximation will also hold for other systems with similar ratios between noise parameters, by scaling the duration of the noise channel accordingly. Modeling other coupling terms can be more subtle, as in the case of an XY coupling term, which we briefly consider in App. H. We show that the composite-channel approximation holds up to time scales for which the non-commutativity of the individual channels becomes nonnegligible. This understanding would be an important a guideline when modelling more general setups than the one considered here, e.g., when incorporating the dynamics during gates, going beyond the code-capacity assumption.
As also discussed in Sec. II, in some quantum devices the qubit parameters vary from qubit to qubit. The natural question that arises is whether the inhomogeneity in the dynamical parameters would have a significant effect on the conclusions of our study. Although checking this question very broadly is beyond the scope of the current work, in App. B we have considered in detail superconducting qubits, and we found that the error-correction dynamics of the spatially inhomogeneous system are still qualitatively similar to the dynamics observed with uniform qubits. We analyzed the non-symmetric reference frame used in current IBM Quantum deployed hardware, the different ways the Pauli approximation can be performed in this case, and the affect it has on the performance of the approximation.
@@
When analyzing the properties of a code, simplified 1Q Pauli noise channels are often assumed in order to facilitate the standard analysis avoiding simulating the logical state. Although for 1Q noise the Pauli approximation holds, we did not manage to find a 1Q approximation that captures the success rate of the code under 2Q ZZ crosstalk, even for very short times. Studying the 2Q Pauli approximations, which could also be used in the standard analysis (without simulating the logical state), we find that its applicability is limited to the low-noise regime (short time scales) which is the most important regime for error correction. It is not a controlled approximation since it is sensitive to the state and decoding scheme, and should be verified carefully with other noise models.
Within the code-capacity assumption, examining the pseudo-threshold in its restricted definition used in this work, we see that it is pretty similar for all approximations, for parameters of low and comparable crosstalk and detuning errors. For devices with such parameters, e.g., current Heron devices, for times shorter than the pseudo-threshold, both the Pauli and the composite-channel approximations worked well. Thus our analysis is a first step towards a detailed analysis of logical error rates under a model that is relatively close to the real hardware, with more elaborate codes.
We have also considered the adaptation of the decoder to the device connectivity with dominant 2Q errors. The ZZ crosstalk noise term is not merely biased noise or correlated noise, but rather it has all the properties of being biased, correlated, and coherent, and does not commute with the amplitude-damping channel. Our computation accounting for all those effects together allowed us to explore the parameter regimes where the modified decoder is advantageous, and generalizing and extending this approach could hold promise for future hardware-tailored codes and decoding schemes.
An interesting property of error-correction codes is that the syndrome measurement changes the noise channel, and some noise terms disappear altogether. Many of the non-Pauli terms disappear this way, although not all of them, as explained in App. F. It raises the question of whether this observation can be used to design an approximation that will be both easy to simulate and accurate, by keeping the most relevant of the non-vanishing noise terms.
@@
Finally, a fundamental open question is how do our results scale with the number of qubits in a code, and how properties of families of codes are affected. In particular, how the error correction threshold scales and changes in the different approximations. This direction could be nontrivial to pursue since it would require the simulation of the error correction procedure with large codes. This may be difficult with a dynamical simulation, but perhaps feasible with the composite-channel approximation, and in particular, using tensor-network methods that are suitable for the relevant states [19]. An understanding of this question may be useful to shed a light on the feasibility of error correction under realistic noise and could possibly contribute to its realization.
@@ Appendix C Pseudo-threshold computations
Notice that we compared the same noise channel on the physical qubit and on the logical qubit. Remembering that the physical qubit is much simpler than the logical qubit, we consider also the case where simulating the dynamical simulation of the noise is possible for the physical qubit, and not for the logical qubit. So when comparing to the curve of the dynamical simulation of the noise on the physical qubit, the pseudo-thresholds of the composite approximations do not change, but the Pauli noise pseduo-threshold becomes $5.84\,\mu$s. This means that only a small part of the mismatch was caused by the inaccuracy of the Pauli approximation on the physical qubit.
@@ I.2 Failure for ZZ crosstalk and large damping noise
Appendix K Modified Decoder details

### v2

@@ V.7 Nonsymmetric reference frame
However, when we are interested in the Pauli approximation, these two representations for the nonsymmetric reference frame are different. This stems from the fact that the Pauli approximation is done on the 1Q noise terms and 2Q noise terms separately. These differences can be seen in Fig. 7, for $\tilde{h}=2\pi\times 5\,$kHz, ${\zeta}=2\pi\times 3\,$kHz, $T_{1}=150$ $\mu$s and $T_{2}=100$ $\mu$s. In Fig. 7(a) we see that for the initial state $|+\rangle_{L}$ in the “2Q-adjusted” representation, the Pauli approximation begins to deviate already at $t=0.3$ $\mu$s, while in the “1Q-adjusted” representation, the Pauli approximation remains accurate up to longer times. The initial state $|0\rangle_{L}$, on the other hand is significantly different, in that the Pauli approximation does not perform well regardless of the representation, as can be seen in Fig. 7(b). In total, it seems that for these noise parameters, the 1Q-adjusted representation is somewhat beneficial for the Pauli approximation, although again, we see that the Pauli approximation is sensitive to the specific noise and state parameters.
V.8 Dynamics of syndrome cycles
In this subsection we go beyond the picture studied so far where the failure rate of codes has been presented for a noise channel acting for varying durations on the system, following which an ideal procedure of syndrome extraction and recovery has been assumed. We also depart in this section from the quantity $\eta$ that has been our measure of the recovery failure rate so far, and we consider matrix elements of the logical density matrix following repeated error correction syndrome cycles.
We employ the modified FC ZZ decoder and we restrict the noise terms to those compatible with this decoder,
$h=2\pi\times 5\,{\rm kHz},\qquad\zeta=2\pi\times 8\,{\rm kHz},$ (47)
and for the dissipative parameters,
$T_{1}=\infty\,\mu{\rm s},\qquad T_{2}=200\,\mu{\rm s}.$ (48)
[…]
Nevertheless, this heuristic factor is not sufficient to fix the Pauli approximation and as the figure shows, after about 200 cycles the state in the dynamical simulation begins to deviate significantly from the scaled Pauli approximation, which significantly underestimates the error (and remains constantly below it). Notably, the value of $\alpha$ appears to oscillate towards a possible steady state (which is seen to be 1/2), and the coherence decays. It is plausible that the coherent phase accumulating within the logical subspace (that is not modeled at all within the Pauli approximation) contributes to this different dynamical evolution. The existence of such rich dynamics on the scale of a large number of syndrome cycles cannot be captured within the stochastic Pauli model, which carries no coherence between cycles.
@@ VI Summary and Outlook
This work is focused on the importance of physically-motivated noise dynamics when determining the success rate of an error correction code to recover logical states subject to noisy dynamics. The Lindblad model studied in this work includes the dominant idle errors in superconducting quantum computing systems [20], or can be modified in order to better model those  [^5] We note that the effect of frequency shifts caused by stochastic charge-parity fluctuation in superconducting qubits [20], can be absorbed into each qubit’s $h_{i}$ value and averaged over the possible frequencies., and could be relevant for other systems as well [64, 65, 66, 67, 68, 69]. As mentioned in Sec. II, in some quantum devices the qubit parameters vary from qubit to qubit. The natural question that arises is how an inhomogeneity in the dynamical parameters would affect the conclusions of our study. In App. B we have considered the five-qubit code dynamics with some spread in the parameters (relevant to superconducting qubits), and we see that the error-correction dynamics of the spatially inhomogeneous system could still be qualitatively similar to the dynamics observed with uniform qubits.
@@
Overall, we have determined that for a Lindbladian of the form presented in Sec. II, the composite-channel approximation (constructed explicitly to account for 2Q noise) is a very good approximation for relevant noise parameters and the time scales that we study. We show that the composite-channel approximation holds up to time scales for which the non-commutativity of the individual channels becomes nonnegligible. This approximation will also hold for other systems with similar ratios between noise parameters, by scaling the duration of the noise channel accordingly. Modeling other coupling terms can be more subtle due to noncommuting terms possibly contributing in shorter time scales, as in the case of an XY coupling term, which we briefly consider in App. H. This understanding would be an important a guideline when modelling more general setups than the one considered here, e.g., when incorporating the dynamics during gates involving parity-check qubits. We have also considered the effect of choosing a reference frame, such as the non-symmetric reference frame often used with superconducting qubits hardware, the different ways the Pauli approximation can be performed in this case, and the affect it has on the performance of the approximation.
@@
When analyzing the properties of a code, simplified 1Q Pauli noise channels are often assumed in order to facilitate the standard analysis. Although for 1Q noise the Pauli approximation holds, we did not manage to find a 1Q approximation that captures the success rate of the code under 2Q ZZ crosstalk, even for very short times, as explained in App. J. Studying the 2Q Pauli approximations we find that its applicability is limited to the low-noise regime (short time scales), which is the most important regime for error correction. However, it is not a controlled approximation since it can be sensitive to the details of the combination of the state and decoding scheme (in addition to interactions between noise terms), and we pointed out significant mechanisms of its failure. We saw several scenarios where the Pauli noise underestimates the logical error rate and hence cannot be taken as a pessimistic model for the error correction rates.
Beyond the feasibility of simulating large code families, the Pauli approximation is prevalent in error correction studies because an essential property of quantum error correction is that the syndrome measurement changes the noise channel, and many of the non-Pauli terms disappear this way, as discussed in App. F. However, since some non-Pauli terms still survive even after the syndrome measurement, this approximation is not exact. In fact, we have seen that after a large number of syndrome cycles, the success rate predicted by the Pauli approximation as we have constructed it misses the underlying dynamics and there does not seem to be an immediate way to remedy this since the culprit seems to be the accumulation of small phase contributions that are inaccessible with Pauli channels. Accounting for imperfect syndrome cycles using a circuit model could be expected to introduce new complexities – e.g., the accumulation of a coherent phase between states in the code space and outside of it, which was not possible in the current model.
@@
Therefore, a fundamental open question is how our observations would change with circuit-based non-Pauli dynamics of error correction, and what would be the relevant effects with larger codes with an increasing distance, and some randomness in the parameters. In particular, how the error correction threshold and logical error rates scale and change when treated within the different approximations. This direction could be nontrivial to pursue since it would require the simulation of the error correction procedure with large codes. This may be difficult with a dynamical simulation, but perhaps feasible with the composite-channel approximation, and in particular, using tensor-network methods that are suitable for the relevant states [20]. An understanding of this question may be useful to shed a light on the feasibility of error correction under realistic noise and could possibly contribute to its realization.
@@ Appendix C Pseudo-threshold computations
C.1 Pseudo-threshold computations
We are presenting in Fig. 11 an example of the plots from which we extract the pseduo-threshold as the crossing point between the $\eta$ of the logical qubit and the infidelity of the physical qubit.
Figure 11: The identification of the pseudo-threshold as the crossing point between the $\eta$ of the logical qubit and the infidelity of the physical qubit, for noise parameters as in Fig. 6. (a) for the dynamical noise (b) for the Pauli noise
Figure 12: The failure probability after error recovery – $\eta$ of Eq. (11) – vs. $t$, the duration of the noise channel, and the noise parameters in Eqs. (40)-(41), with dominant crosstalk noise on an average over the initial states in Eq. (9), in the short-time regime, showing that all approximations are identical. The pseudo-threshold for this case is $t=0.069\mu s$ and we can see that all approximations are identical up to the pseudo-threshold.
C.2 Pseudo-threshold for other noise parameters
When computing the physical error rate, we assumed a single qubit standing idle, and so it does not suffer from two-qubit interaction at all. In a sense this gives an advantage to the physical qubit, since it is oblivious to an important noise source. In practice, when the qubits interact, the physical qubit will also suffer from two-qubit noise, but since we are analyzing only a single logical qubit, we are not dealing with that for now. And so, in the case of pure ZZ noise, where there is no single qubit noise at all, then the physical qubit does not suffer from any noise, and so there will be no pseudo-threshold. Also, for the cases with eagle devices the threshold will be extremely low, because the dominant noise source is the two qubit interaction. For completeness, we present in Fig. 12 the $\eta$ for the eagle parameters as in Fig. 2 with the pseudo-threshold.
C.3 Dynamical noise on a physical qubit
We compared the same noise channel on the physical qubit and on the logical qubit. Remembering that the physical qubit is much simpler than the logical qubit, we consider also the case where simulating the dynamical simulation of the noise is possible for the physical qubit, and not for the logical qubit. So when comparing to the curve of the dynamical simulation of the noise on the physical qubit, the pseudo-thresholds of the composite approximations do not change, but the Pauli noise pseudo-threshold becomes $5.84\,\mu$s. This means that only a small part of the mismatch was caused by the inaccuracy of the Pauli approximation on the physical qubit.
@@ I.2 Failure for ZZ crosstalk and large damping noise
Appendix K A logical error with the Pauli crosstalk channel
In this subsection we point to the reason that the Pauli channel constructed from the 2Q crosstalk generators of the connected pairs could deviate from the unitary dynamics that it is supposed to describe, when the full connectivity ZZ decoder is applied, to some input states.
For this purpose we define the crosstalk coefficients
$c=\cos\frac{\zeta t}{2},\qquad s=\sin\frac{\zeta t}{2}.$ (84)
The action of a Pauli noise channel composed of $n$ compositions of the 2Q Pauli channel defined in Eq. (34) can be written in terms of a vectorized superoperator as $\hat{P}\ket{\rho}\!\rangle$, and $\hat{P}$ expanded to second order terms as
$\hat{P}=c^{2n}\hat{1\kern-2.5pt\text{l}}+c^{2(n-1)}s^{2}\sum_{\langle i,j\rangle}\hat{Z}_{i}\hat{Z}_{j}+\\ c^{2(n-2)}s^{4}\sum_{\langle i,j\rangle}\hat{Z}_{i}\hat{Z}_{j}\sum_{\langle k,l\rangle\neq\langle i,j\rangle}\hat{Z}_{k}\hat{Z}_{l}+...$ (85)
where the summation is over the connected pairs and it is understood that the hat notation of each of the terms in the above expansion indicates that it acts simultaneously on the left and on the right of the density matrix (as discussed in App. E).
The unitary operator constructed from the application of the same generators would be
$U=\exp\left\{-i\sum_{\langle i,j\rangle}\frac{\zeta t}{2}Z_{i}Z_{j}\right\}=\prod_{\langle i,j\rangle}\left(c1\kern-2.5pt\text{l}-isZ_{i}Z_{j}\right),$ (86)
which can be expanded to the same order of terms,
$U=u_{0}+u_{1}+u_{2}+...$ (87)
with
$u_{0}=c^{n}1\kern-2.5pt\text{l},\qquad u_{1}=(-i)c^{n-1}s\sum_{\langle i,j\rangle}Z_{i}Z_{j},$ (88)
and
$u_{2}=(-1)c^{n-2}s^{2}\sum_{\langle i,j\rangle}Z_{i}Z_{j}\sum_{\langle k,l\rangle\neq\langle i,j\rangle}Z_{k}Z_{l}.$ (89)
The action of $U$ is given by $\rho\to U\rho U^{\dagger}$, and we are interested in spotting the leading-order difference from the action of $\hat{P}$ that gives an error not correctable by the code. Since the code corrects any $ZZ$ error, the leading order difference would come from a combination of four different (non-cancelling) $Z$ factors, i.e.,
$Z_{i}Z_{j}Z_{k}Z_{l},\quad i\neq j\neq k\neq l,$ (90)
which is equivalent to the logical $Z_{L}$ up to the missing qubit’s $Z_{m}$, and could result in $Z_{L}$ being applied to the state if the syndrome would be interpreted (and corrected) as the single-qubit error $Z_{m}$. When the input state is not an eigenvalue of the logical $Z_{L}$, this results in a logical error.
Terms as in Eq. (90) appear in the expansion of $U\rho U^{\dagger}$ in leading order in the term $u_{2}\rho u_{2}^{\dagger}$, which is explicitly,
$c^{2(n-2)}s^{4}\sum_{\langle i,j\rangle}Z_{i}Z_{j}\sum_{\langle k,l\rangle}Z_{k}Z_{l}(\rho)\sum_{\langle i^{\prime},j^{\prime}\rangle}Z_{i^{\prime}}Z_{j^{\prime}}\sum_{\langle k^{\prime},l^{\prime}\rangle}Z_{k^{\prime}}Z_{l^{\prime}},$ (91)
with $\langle i,j\rangle\neq\langle k,l\rangle$ and $\langle i^{\prime},j^{\prime}\rangle\neq\langle k^{\prime},l^{\prime}\rangle$. The above expression should be contrasted with the Pauli construction of $\hat{P}$ that gives at the same order of terms (with $\langle i,j\rangle\neq\langle k,l\rangle$ as before),
$c^{2(n-2)}s^{4}\sum_{\langle i,j\rangle}Z_{i}Z_{j}\sum_{\langle k,l\rangle}Z_{k}Z_{l}(\rho)Z_{i}Z_{j}Z_{k}Z_{l},$ (92)
and a simple combinatorial counting shows that the expression in Eq. (91) contains three times more terms that do not cancel and lead to a logical error larger by a factor of three as compared with Eq. (92). This is easiest to see by fixing one term on the left of $\rho$ in Eq. (91), e.g., $(i,j,k,l)=(1,2,3,4)$, which can be matched by three terms to the right of $\rho$, i.e.,
$(i^{\prime},j^{\prime},k^{\prime},l^{\prime})=(1,2,3,4),\,(1,3,2,4),\,(1,4,2,3),$ (93)
of which only the first exists in the Pauli construction. This results in a factor of 3 that in this case – due to the $s^{4}$ dependence if Eq. (92) – can be corrected for small $|\zeta|t$ by a scaling of the noise by $\zeta\to\sqrt[4]{3}\zeta$ to result in the same logical error (up to higher orders in $\zeta$).
The above expansion could perhaps be generalized numerically or circumvented with some scaling factors when building the Pauli noise, in order to accommodate for qubit-dependent phase terms, and other types of noise and qubit connectivities.
Appendix L Modified Decoder details